# Data-driven Model Predictive Control: Asymptotic Stability despite Approximation Errors exemplified in the Koopman framework

### v1

@@ 3. Asymptotic stability of MPC
In the following we derive upper bounds for the two terms appearing in (18), that are $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|$ and $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|$. Consider $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|$. For $i=0$, we use the proportional error bound of Assumption 1 to derive that
$\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(0;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(0;x^{+})\|=\|\tilde{x}^{+}-x^{+}\|=\|f^{\varepsilon}(\hat{x},\mu_{N}^{\varepsilon}(\hat{x}))-f(\hat{x},\mu_{N}^{\varepsilon}(\hat{x}))\|\leq c_{x}^{\varepsilon}\|\hat{x}\|+c_{u}^{\varepsilon}\|\mu_{N}^{\varepsilon}(\hat{x})\|.$
For $i\geq 1$ we use that, in view of Assumption 3, $f^{\varepsilon}$ is Lipschitz continuous with Lipschitz constant $L_{f^{\varepsilon}}$, to obtain
$\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|\leq L_{f^{\varepsilon}}^{i}\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(0;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(0;x^{+})\|\leq L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}\|\hat{x}\|+c_{u}^{\varepsilon}\|\mu_{N}^{\varepsilon}(\hat{x})\|).$
Consider $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|$. For the growth bound of $f^{\varepsilon}$ derived in Proposition 1 and for the relaxed Lyapunov inequality (15), we have that
$V_{N}^{\varepsilon}(f^{\varepsilon}(\hat{x},\mu_{N}^{\varepsilon}(\hat{x})))$ $=\sum_{i=0}^{N-1}\ell(x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+}),u^{\sharp}(i))$
@@
we also have
$\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|_{Q}^{2}\leq B^{\varepsilon}_{N}\|\hat{x}\|_{Q}^{2}.$
By exploiting standard inequalities of weighted squared norms and taking the square root, we have that
$\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|\leq\sqrt{B^{\varepsilon}_{N}\frac{\bar{\lambda}}{\underline{\lambda}}}\|\hat{x}\|.$
Then, substituting the previous inequalities in (18), we obtain
@@
$\hskip 14.22636pt\leq 2\|Q\|\sqrt{B^{\varepsilon}_{N}\frac{\bar{\lambda}}{\underline{\lambda}}}\|\hat{x}\|L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}\|\hat{x}\|+c_{u}^{\varepsilon}\|\mu_{N}^{\varepsilon}(\hat{x})\|)+\|Q\|L_{f^{\varepsilon}}^{2i}(c_{x}^{\varepsilon}\|\hat{x}\|+c_{u}^{\varepsilon}\|\mu_{N}^{\varepsilon}(\hat{x})\|)^{2}$
$\hskip 14.22636pt\leq 2\|Q\|\sqrt{B^{\varepsilon}_{N}\frac{\bar{\lambda}}{\underline{\lambda}}}L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}\|\hat{x}\|^{2}+\underbrace{c_{u}^{\varepsilon}\|\hat{x}\|\|\mu_{N}^{\varepsilon}(\hat{x})\|}_{\leq\frac{1}{2}c_{u}^{\varepsilon}(\|\hat{x}\|^{2}+\|\mu_{N}^{\varepsilon}(\hat{x})\|^{2})})+2\|Q\|L_{f^{\varepsilon}}^{2i}\Big{(}(c_{x}^{\varepsilon})^{2}\|\hat{x}\|^{2}+(c_{u}^{\varepsilon})^{2}\|\mu_{N}^{\varepsilon}(\hat{x})\|^{2}\Big{)}.$
Substituting this back in (17) and (16), we obtain that
@@
$C_{x}$ and $C_{u}$ can be made arbitrarily small with a sufficiently small proportionality constants $c_{x}^{\varepsilon}$ and $c_{u}^{\varepsilon}$. Then there exist $\bar{\alpha}\in(0,1)$ and $\bar{c}_{x},\bar{c}_{u}>0$ such that
$-\alpha^{\varepsilon}\ell(\hat{x},\mu_{N}^{\varepsilon}(\hat{x}))+C_{x}\|\hat{x}\|^{2}+C_{u}\|\mu_{N}^{\varepsilon}(\hat{x})\|^{2}<-\bar{\alpha}\|\hat{x}\|^{2}$
for all $c_{x}^{\varepsilon}\leq\bar{c}_{x}$ and $c_{u}^{\varepsilon}\leq\bar{c}_{u}$, proving asymptotic stability of the closed-loop. $\square$
Based on Theorem 1 we can now even prove cost controllability of the surrogate model, i.e., existence of a bounded sequence $(\tilde{B}^{\varepsilon}_{k})_{k\in\mathbb{N}}$ satisfying the growth bound (8) on the sub-level set $V_{c}^{-1}$ if we modify the set of admissible sets based on the knowledge of a sufficiently large prediction horizon $N$, for which we were able to conclude asymptotic stability in Theorem 3. To be more precise, we have to replace the set $\mathcal{B}_{k\eta}(0)$ by $\mathcal{B}_{\min\{N,k\}\eta}(0)$ in (7) to ensure that the set of admissible controls remains nontrivial.
@@
Let the assumptions of Theorem 1 hold. Then there exists a monotonically increasing and bounded sequence $(\tilde{B}_{k}^{\varepsilon})_{k\in\mathbb{N}}$ such that cost controllability holds for the surrogate model, i.e. for all $\hat{x}\in V_{c}^{-1}$ there exists $\mathbf{u}\in\tilde{\mathcal{U}}^{\varepsilon}_{\infty}(\hat{x})$ such that
$V^{\varepsilon}_{\bar{N}}(\hat{x})\leq J^{\varepsilon}_{\bar{N}}(\hat{x},\mathbf{u})\leq\tilde{B}^{\varepsilon}_{\bar{N}}\ell^{\star}(\hat{x})\qquad\forall\,\bar{N}\in\mathbb{N}$
where $\tilde{\mathcal{U}}^{\varepsilon}_{\infty}(\hat{x})$ is defined as in (7) replacing the set $\mathcal{B}_{k\eta}(0)$ by $\mathcal{B}_{\min\{N,k\}\eta}(0)$.
Proof. The proof resembles the proof of Proposition 6 in MullWort17 . In the proof of Theorem 1, in (15), we showed that there exists a sufficiently small $\varepsilon$ such that
$V^{\varepsilon}_{N}(f^{\varepsilon}(x,\mu^{\varepsilon}_{N}(x)))\leq V^{\varepsilon}_{N}(x)-\alpha^{\varepsilon}\ell(x,\mu^{\varepsilon}_{N}(x)).$
Then, denoting by $\mathbf{u}_{\mathrm{MPC}}$ the input sequence obtained applying the MPC control law $\mu^{\varepsilon}_{N}$, we can relate the infinite-horizon optimal cost to $\alpha^{\varepsilon}$
$V_{\infty}^{\varepsilon}(\hat{x})$ $\leq J_{\infty}^{\varepsilon}(\hat{x},\mathbf{u}_{\mathrm{MPC}})\leq\sum_{i=0}^{\infty}\ell(x^{\varepsilon}(i),\mu_{N}^{\varepsilon}(x^{\varepsilon}(i)))$
$\leq(\alpha^{\varepsilon})^{-1}\sum_{i=0}^{\infty}V_{N}^{\varepsilon}(x^{\varepsilon}(i))-V_{N}^{\varepsilon}(f(x^{\varepsilon}(i),\mu_{N}^{\varepsilon}(x^{\varepsilon}(i)))$
$\leq(\alpha^{\varepsilon})^{-1}V_{N}^{\varepsilon}(\hat{x})\leq(\alpha^{\varepsilon})^{-1}B_{N}^{\varepsilon}\ell^{\star}(\hat{x}).$
Since $V_{N}^{\varepsilon}(\hat{x})\leq V_{N+1}^{\varepsilon}(\hat{x})\leq V_{\infty}^{\varepsilon}(\hat{x})$ for all $N\in\mathbb{N}$, it follows that
$V_{N}^{\varepsilon}(\hat{x})\leq(\alpha^{\varepsilon})^{-1}B_{N}^{\varepsilon}\ell^{\star}(\hat{x}).$
Hence, the statement holds for
$\tilde{B}_{k}^{\varepsilon}:=\min\{B_{k}^{\varepsilon},(\alpha^{\varepsilon})^{-1}B_{N}^{\varepsilon}\}.$
$\square$
@@ 4. Data-driven models via kernel EDMD
In this section, we show how the assumptions used to derive asymptotic stability of the MPC closed-loop in Section 3 can be met using Koompan-based surrogate models. To this end, we leverage kernel extended dynamic mode decomposition (kEDMD) and its extension to control systems proposed in BoldPhil24 as a data-driven method to approximate the Koopman operator.
Let $\mathsf{k}:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a continuous, symmetric and strictly positive-definite kernel function, i.e., for every set $\mathcal{Z}=\{z_{1},\dots,z_{p}\}\subset\mathbb{R}^{n}$ of pairwise distinct elements, the kernel matrix $K_{\mathcal{Z}}=(\mathsf{k}(z_{i},z_{j}))_{i,j=1}^{p}$ is positive definite. For $z\in\mathbb{R}^{n}$, the canonical features $\Phi_{z}$ of $\mathsf{k}$ are defined by $\Phi_{z}(x)=\mathsf{k}(z,x)$, $x\in\mathbb{R}^{n}$. By completion, the kernel $\mathsf{k}$ induces an Hilbert space $\mathbb{H}$ with inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$, see Wend04 . Importantly, elements $f\in\mathbb{H}$ fulfill the reproducing property
$f(x)=\langle f,\Phi_{x}\rangle_{\mathbb{H}}\qquad\forall\,x\in\mathbb{R}^{n}.$
Hence, $\mathbb{H}$ is called reproducing kernel Hilbert space (RKHS).
Remark 1 (Wendland kernels).
In this work, we use piecewise-polynomial and compactly-supported kernel functions based on the Wendland radial basis functions (RBF) $\Phi_{n,k}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ with smoothness degree $k\in\mathbb{N}$. The RKHS induced by the Wendland kernels coincides with fractional Sobolev spaces with equivalent norms Wend04 , a key property, which also holds for Matérn kernels; see FassYe11 . The induced (Wendland) kernel is given by
$\mathsf{k}(x,y)=\phi_{n,k}(\|x-y\|)\qquad\text{for }x,y\in\mathbb{R}^{n},$
see (Wend04, , Table 9.1). In the numerical simulations of Section 5, $k=1$ is used, which corresponds to $\phi_{n,1}\in\mathcal{C}^{2}([0,\infty),\mathbb{R})$ defined by $\phi_{n,1}(r)=\frac{1}{20}(1-r)^{4}(4r+1)$ for $r<1$ (and 00 otherwise) for $n\leq 3$ and by $\phi_{n,1}(r)=\frac{1}{30}(1-r)^{5}(5r+1)$ for $r<1$ (and 00 otherwise) for $n\in\{4,5\}$.
@@ 4.1. Koopman operator and EDMD
Let $\Omega\subset\mathbb{R}^{n}$ be an open and bounded set with Lipschitz boundary containing the origin $0=0_{\mathbb{R}^{n}}$ in its interior. For a map $F:\Omega\rightarrow\mathbb{R}^{n}$, we consider the autonomous discrete-time dynamical system given by
$x^{+}=F(x).$ (DS)
In the following, we assume that the set $\Omega$ is forward invariant w.r.t. the dynamics $F$ of (DS), i.e., $F(x)\in\Omega$ for all $x\in\Omega$, to streamline the presentation and refer to kohne2024infty for the necessary (technical) modifications if this assumption does not hold. The associated (infinite-dimensional) linear and bounded Koopman operator $\mathcal{K}:\mathcal{C}_{b}(\Omega)\rightarrow\mathcal{C}_{b}(\Omega)$ is defined by the identity
@@
Let $\mathcal{K}|_{V_{\mathcal{X}}}$ be the restriction of the Koopman operator on the subspace $V_{\mathcal{X}}$. As introduced in kohne2024infty , we consider the matrix approximant $\widehat{K}$ of $\mathcal{K}|_{V_{\mathcal{X}}}$, which is given by
@@
where $K_{F(\mathcal{X})}=(\mathsf{k}(x_{i},F(x_{j})))_{i,j=1}^{d}$. In the following, we tacitly associate $\widehat{K}$ with the induced linear map from $V_{\mathcal{X}}$ to $V_{\mathcal{X}}$. The approximation $\widehat{K}$ corresponds to the widely used kernel extended dynamic mode decomposition (kEDMD; williams2014kernel ). For an observable function $\psi\in\mathbb{H}\subseteq\mathcal{C}_{b}(\Omega)$, the surrogate dynamics are given by
$\psi(F(x))\approx\psi^{+}(x):=\sum_{i=1}^{d}(\widehat{K}\psi_{\mathcal{X}})_{i}\Phi_{x_{i}}(x),$
where $\psi_{\mathcal{X}}=(\psi(x_{1}),\dots,\psi(x_{d}))^{\top}$. A bound on the full approximation error was established in (kohne2024infty, , Theorem 5.2) and is provided in the following theorem. Therein, the approximation accuracy of kEDMD can be described using the fill distance $h_{\mathcal{X}}$ of the set $\mathcal{X}$ in $\Omega$, which is defined by
@@
Let $\mathbb{H}$ be the RKHS on $\Omega$ generated by the Wendland kernels with smoothness degree $k\in\mathbb{N}$. Let the right-hand side ${F\in\mathcal{C}^{p}_{b}(\Omega,\mathbb{R}^{n})}$ of system (DS), with $p=\lceil\frac{n+1}{2}+k\rceil$. Then, there exist constants $C,h_{0}>0$ such that the bound on the full approximation error
@@
holds for all sets $\mathcal{X}:=\{x_{i}\mid i\in[1:d]\}\subset\Omega$, $d\in\mathbb{N}$, of pairwise-distinct data points with fill distance $h_{\mathcal{X}}$, $h_{\mathcal{X}}\leq h_{0}$.
We note that, for a large number of data points in $\mathcal{X}$, the computation of the approximation of the Koopman operator (20) can be numerically instable due to bad conditioning of the kernel matrix $K_{\mathcal{X}}$. To alleviate this, one may include regularization by replacing ${K}_{\mathcal{X}}^{-1}$ with $({K}_{\mathcal{X}}+\lambda I)^{-1}$ with regularization parameter $\lambda>0$. Error bounds in the spirit of Theorem 2 for this regularized surrogate were proven in (BoldPhil24, , Thm. 2.4).

### v2

@@ 3. Asymptotic stability of MPC
[…]
First, we consider the term $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|$. For $i=0$, we leverage (P-bound) of Assumption 1 to infer
$\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(0;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(0;x^{+})\|=\|\tilde{x}^{+}-x^{+}\|$
[…]
Using Lipschitz continuity of the surrogate model (3) given by $f^{\varepsilon}$ with Lipschitz constant $L_{f^{\varepsilon}}$ due to Assumption 2, we obtain for $i\geq 1$
$\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;x^{+})\|$ $\leq$ $L_{f^{\varepsilon}}^{i}\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(0;\tilde{x}^{+})-x_{\mathbf{u}^{\sharp}}^{\varepsilon}(0;x^{+})\|$
$\leq$ $).\displaystyle L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}\|\hat{x}\|+c_{u}^{\varepsilon}\|\mu_{N}^{\varepsilon}(\hat{x}$$\|)$.
[…]
@@
we also have $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|_{Q}^{2}\leq B^{\varepsilon}_{N}\|\hat{x}\|_{Q}^{2}$. By exploiting standard inequalities of weighted squared norms and taking the square root, we have $\|x_{\mathbf{u}^{\sharp}}^{\varepsilon}(i;\tilde{x}^{+})\|\leq\sqrt{B^{\varepsilon}_{N}\bar{\lambda}/\underline{\lambda}}\|\hat{x}\|$. Then, substituting this inequalities in (15), we get
@@
$\leq$ $)\displaystyle 2\|Q\|\sqrt{B^{\varepsilon}_{N}\frac{\bar{\lambda}}{\underline{\lambda}}}\|\hat{x}\|L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}\|\hat{x}\|+c_{u}^{\varepsilon}\|\mu_{N}^{\varepsilon}(\hat{x}$$\|)$
$+\|Q\|L_{f^{\varepsilon}}^{2i}(c_{x}^{\varepsilon}\|\hat{x}\|+c_{u}^{\varepsilon}\|\mu_{N}^{\varepsilon}(\hat{x})\|)^{2}$
$\leq$ $2\|Q\|\sqrt{B^{\varepsilon}_{N}\frac{\bar{\lambda}}{\underline{\lambda}}}L_{f^{\varepsilon}}^{i}(c_{x}^{\varepsilon}\|\hat{x}\|^{2}+\underbrace{c_{u}^{\varepsilon}\|\hat{x}\|\|\mu_{N}^{\varepsilon}(\hat{x})\|}_{\leq\frac{1}{2}c_{u}^{\varepsilon}(\|\hat{x}\|^{2}+\|\mu_{N}^{\varepsilon}(\hat{x})\|^{2})})$
$+2\|Q\|L_{f^{\varepsilon}}^{2i}\Big((c_{x}^{\varepsilon})^{2}\|\hat{x}\|^{2}+(c_{u}^{\varepsilon})^{2}\|\mu_{N}^{\varepsilon}(\hat{x})\|^{2}\Big).$
Substituting this back in (14) and (13), we obtain
@@
$C_{x}$ and $C_{u}$ can be made arbitrarily small with sufficiently small proportionality constants $c_{x}^{\varepsilon}$ and $c_{u}^{\varepsilon}$. Then, there exists a sufficiently small $\varepsilon_{0}$ (potentially further tightened in comparison to our first choice resulting from the computation of the lower bound $\alpha_{N}^{\varepsilon_{0}}$ on the suboptimality index) such that $\bar{c}_{x}$, $\bar{c}_{u}$ are sufficiently small to ensure the inequality
$\alpha^{\varepsilon}_{N}\ell(\hat{x},\mu_{N}^{\varepsilon}(\hat{x}))+C_{x}\|\hat{x}\|^{2}+C_{u}\|\mu_{N}^{\varepsilon}(\hat{x})\|^{2}<-\bar{\alpha}\|\hat{x}\|^{2}$
for some $\bar{\alpha}\in(0,\alpha_{N}^{\varepsilon_{0}})$ and for all $\varepsilon\in(0,\varepsilon_{0})$. This completes the proof and, thus, shows asymptotic stability of the MPC closed loop based on the surrogate model (3). $\square$
The theoretical guarantees provided by Theorem 1 should be interpreted in a qualitative way, i.e., asymptotic stability is preserved provided that the approximation is sufficiently accurate and the optimization horizon $N$ sufficiently large. However, the derived formulas for the minimum horizon and the maximum error bounds will, in general, be rather conservative in practice.
Based on Theorem 1 we can now even prove cost controllability of the surrogate model, i.e., existence of a bounded sequence $(B^{\varepsilon}_{k})_{k\in\mathbb{N}}$ satisfying the growth bound (5) on the sublevel set $V_{N}^{-\varepsilon}(c)$ of Theorem 1.
@@
Let the assumptions of Theorem 1 hold. Then there exists a monotonically increasing and bounded sequence $(B_{k}^{\varepsilon})_{k\in\mathbb{N}}$ such that cost controllability holds for the surrogate model on the set $V_{N}^{-\varepsilon}(c)\subseteq S$, i.e. for all pairs $(\hat{x},N)\in V_{N}^{-\varepsilon}(c)\times\mathbb{N}$, there exists $\mathbf{u}\in\mathcal{U}_{N}$ satisfying
[…]
Proof. The proof resembles the proof of Proposition 6 in Müller and Worthmann (2017). In the chain of inequalities (12), see the proof of Theorem 1, we showed that there exists a sufficiently small $\varepsilon$ such that
$V^{\varepsilon}_{N}(f^{\varepsilon}(x,\mu^{\varepsilon}_{N}(x)))\leq V^{\varepsilon}_{N}(x)-\alpha^{\varepsilon}_{N}\ell(x,\mu^{\varepsilon}_{N}(x))$
holds for all $x\in V_{N}^{-\varepsilon}(c)$. Then, denoting by $\mathbf{u}_{\mathrm{MPC}}$ the input sequence obtained applying the MPC control law $\mu^{\varepsilon}_{N}$, we can relate the infinite-horizon optimal cost to $\alpha^{\varepsilon}_{N}$
[…]
$\leq(\alpha^{\varepsilon}_{N})^{-1}\sum_{i=0}^{\infty}V_{N}^{\varepsilon}(x^{\varepsilon}(i))-V_{N}^{\varepsilon}(f(x^{\varepsilon}(i),\mu_{N}^{\varepsilon}(x^{\varepsilon}(i)))$
[…]
Hence, the growth bounds holds with $B_{k}^{\varepsilon}:=B_{N}^{\varepsilon}/\alpha^{\varepsilon}_{N}$, $k>N$, which shows the assertion. $\square$
Corollary 1 shows that cost controllability is inherited by the data-driven surrogate model (3) provided that Assumptions 1 and 2 hold. This result extends Proposition 1, in which the growth bound is only shown for finitely many optimization horizons $N$, $N\in[1:\bar{N}]$.
Remark 1.
We point out that the sets $\Omega$ and $S$ are not employed in the MPC algorithm meaning that we have not yet incorporated state constraints. Rather, we assume cost controllability of the original system (1) on a set $S$ and, then, require Assumptions 1 and 2 on a (sufficiently large) set $\Omega$, such that $S\oplus r(\bar{N})\eta^{\varepsilon}\subseteq\Omega$, in which we have to collect data to generate the surrogate model (3). Clearly, a larger set $S$ (and, thus, also a larger set $\Omega$) allows for a larger sublevel set, which corresponds to the (guaranteed) domain of attraction. The incorporation of state constraints is left for future research and requires additional care to ensure recursive feasibility. This might, e.g., be done based on the techniques proposed in Boccia et al. (2014) or using the recently introduced concept of a constraint horizon Do Nascimento et al. (2025), where only a finite number of predicted state are forced to obey the imposed state constraints. Moreover, we incorporated state constraints in a data-driven MPC with stabilizing terminal conditions in our recent work Schimperna et al. (2025a), where the constraints were tightened based on the Lipschitz constant $\bar{L}$ of the surrogate model. Then, combining both ideas using Assumptions 1 and 2 may resolve problems with so-far missing state constraints.
Finally, we emphasize the symmetry in the derived results meaning that verifying cost controllability of the surrogate model (3) for sufficiently small accuracy parameter $\varepsilon$ would suffice to conclude cost controllability of the original system and, thus, also to ensure asymptotic stability of the MPC closed loop based on the surrogate model, but applied to the original system (1).
@@ 4. Data-driven models via kernel EDMD
In this section, we verify Assumptions 1 and 2, i.e., the assumptions imposed on the surrogate model (3), using Koopman-operator theory. To this end, we leverage kernel extended dynamic mode decomposition (kEDMD) and its extension to control systems proposed in Bold et al. (2025b) as a data-driven method to approximate the Koopman operator. Moreover, we briefly discuss kernel approximations of the original dynamics (1) in Subsection 4.3 to emphasize that using the Koopman operator only serves as one of potentially many options to generate data-driven surrogate models satisfying Assumptions 1 and 2.
Let $\mathsf{k}:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a continuous, symmetric and strictly positive-definite kernel function, i.e., for every set $\mathcal{X}=\{x_{1},\dots,x_{d}\}\subset\mathbb{R}^{n}$ of pairwise distinct elements, the kernel matrix $K_{\mathcal{X}}=(\mathsf{k}(x_{i},x_{j}))_{i,j=1}^{p}$ is positive definite. For $x\in\mathbb{R}^{n}$, the canonical features $\Phi_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ of $\mathsf{k}$ are defined by $\Phi_{x}(x^{\prime})=\mathsf{k}(x,x^{\prime})$, $x^{\prime}\in\mathbb{R}^{n}$. By completion, the kernel $\mathsf{k}$ induces an Hilbert space $\mathbb{H}$ with inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$, see Wendland (2004). Importantly, elements $f\in\mathbb{H}$ fulfill the reproducing property
$f(x)=\langle f,\Phi_{x}\rangle_{\mathbb{H}}\qquad\forall\,x\in\mathbb{R}^{n}$
showing that point evaluation is well defined in $\mathbb{H}$. $\mathbb{H}$ is called reproducing kernel Hilbert space (RKHS).
In this work, we use piecewise-polynomial and compactly-supported kernel functions based on the Wendland radial basis functions (RBFs) $\Phi_{n,k}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ with smoothness degree $k\in\mathbb{N}$. The RKHS induced by the Wendland kernels coincides with fractional Sobolev spaces with equivalent norms Wendland (2004), a key property, which also holds for Matérn kernels; see Fasshauer and Ye (2011). The induced (Wendland) kernel is given by
$\mathsf{k}(x,y)=\phi_{n,k}(\|x-y\|)\qquad\text{for }x,y\in\mathbb{R}^{n}$
In the numerical simulations of Section 5, $k=1$ is used, which corresponds to $\phi_{n,1}\in\mathcal{C}^{2}([0,\infty),\mathbb{R})$ defined by
$\phi_{n,1}(r)=\begin{cases}\frac{1}{20}(1-r)^{4}(4r+1)&\text{for }n\in\{2,3\}\\ \frac{1}{30}(1-r)^{5}(5r+1)&\text{for }n\in\{4,5\}\end{cases}$
for $r<1$ (and 0 otherwise), see also (Wendland, 2004, Table 9.1).
@@ 4.1. Koopman operator and EDMD
Let $\Omega\subset\mathbb{R}^{n}$ be an open and bounded set with Lipschitz boundary containing the origin in its interior. We consider the autonomous discrete-time dynamical system given by
$x^{+}=F(x)$ (DS)
with map $F:\Omega\rightarrow\mathbb{R}^{n}$. In the following, we assume that the set $\Omega$ is forward invariant w.r.t. the dynamics (DS), i.e., $F(x)\in\Omega$ for all $x\in\Omega$, to streamline the presentation and refer to Köhne et al. (2025) for the necessary (technical) modifications if this assumption does not hold. The associated (infinite-dimensional) linear and bounded Koopman operator $\mathcal{K}:\mathcal{C}_{b}(\Omega)\rightarrow\mathcal{C}_{b}(\Omega)$ is defined by the identity
@@
Let $\mathcal{K}|_{V_{\mathcal{X}}}$ be the restriction of the Koopman operator on the subspace $V_{\mathcal{X}}$. As introduced in Köhne et al. (2025), we consider the matrix approximant $\widehat{K}$ of $\mathcal{K}|_{V_{\mathcal{X}}}$, which is given by
@@
where $K_{F(\mathcal{X})}=(\mathsf{k}(x_{i},F(x_{j})))_{i,j=1}^{d}$ and $K_{\mathcal{X}}$ the kernel matrix corresponding to $\mathcal{X}$. In the following, we tacitly associate $\widehat{K}$ with the induced linear map from $V_{\mathcal{X}}$ to $V_{\mathcal{X}}$. The approximation $\widehat{K}$ corresponds to the widely used kernel extended dynamic mode decomposition (kEDMD; Williams et al. (2015b)). For an observable function $\psi\in\mathbb{H}\subseteq\mathcal{C}_{b}(\Omega)$, the surrogate dynamics are given by
$\psi(F(x))\approx\psi^{+}(x):=\sum\nolimits_{i=1}^{d}(\widehat{K}\psi_{\mathcal{X}})_{i}\Phi_{x_{i}}(x),$
where $\psi_{\mathcal{X}}=(\psi(x_{1}),\dots,\psi(x_{d}))^{\top}$. A pointwise bound on the full approximation error was established in (Köhne et al., 2025, Theorem 5.2) and is provided in the following theorem. Therein, the approximation accuracy of kEDMD can be described using the fill distance $h_{\mathcal{X}}$ of the set $\mathcal{X}$ in $\Omega$, i.e.,
@@
Let $\mathbb{H}$ be the RKHS on $\Omega$ generated by the Wendland kernels with smoothness degree $k\in\mathbb{N}$. Let $F\in\mathcal{C}^{p}_{b}(\Omega,\mathbb{R}^{n})$} with $p=\lceil\frac{n+1}{2}+k\rceil$ hold for system (DS). Then, there exist constants $C,h_{0}>0$ such that the bound
@@
on the full approximation error holds for all sets $\mathcal{X}:=\{x_{i}\mid i\in[1:d]\}\subset\Omega$, $d\in\mathbb{N}$, of pairwise-distinct data points with fill distance $h_{\mathcal{X}}$, $h_{\mathcal{X}}\leq h_{0}$.
We note that, for a large number of data points in $\mathcal{X}$, the computation of the approximation of the Koopman operator (17) can be numerically instable due to bad conditioning of the kernel matrix $K_{\mathcal{X}}$. To alleviate this, one may include regularization by replacing ${K}_{\mathcal{X}}^{-1}$ with $({K}_{\mathcal{X}}+\lambda I)^{-1}$ with regularization parameter $\lambda>0$. Error bounds in the spirit of Theorem 2 for this regularized surrogate were proven in (Bold et al., 2025b, Theorem 2.4).